# ACC-Debate: An Actor-Critic Approach to Multi-Agent Debate

### v1

@@ 5.3 Individual Performance
Next, we examine the relative effectiveness of both the actor and the critic. To do this, we train an actor and critic via ACC-Debate. Then, during the debate, we pair the trained actor with an untrained critic and pair the trained critic with an untrained actor. In Table 2, column “Actor” corresponds to the former, while “Critic” corresponds to the latter. On average, the trained actor attains higher accuracy compared to the trained critic, this aligns with intuition as the actor is responsible for providing answers while the critic plays a supporting role. In most cases, the trained actor (paired with an untrained critic) outperforms SFT and DebateGPT. When the trained actor is paired with a trained critic (either ACC-Debate or ACC-Debate+), its performance is further improved.
Llama-3 SoM (4x) DebateGPT Actor Critic ACC-Debate ACC-Debate+ BoolQ .$811_{\pm.007}.$$815_{\pm.005}.$$867_{\pm.003}.$$834_{\pm.005}.$$887_{\pm.005}.$$\mathbf{.894}_{\pm.003}$ MMLU .$635_{\pm.004}.$$654_{\pm.005}.$$651_{\pm.012}.$$65_{\pm.008}.$$644_{\pm.01}.$$\mathbf{.683}_{\pm.012}$ BBH .$514_{\pm.005}.$$551_{\pm.008}.$$583_{\pm.01}.$$55_{\pm.015}.$$\mathbf{.593}_{\pm.006}$$574_{\pm.003}.$ SCIQ .$923_{\pm.002}.$$932_{\pm.001}.$$947_{\pm.002}.$$945_{\pm.001}.$$\mathbf{.952}_{\pm.0}$$948_{\pm.003}.$ ARC .$874_{\pm.001}.$$876_{\pm.002}.$$\mathbf{.885}_{\pm.003}$$866_{\pm.002}.$$881_{\pm.004}.$$869_{\pm.002}.$ Mistral SoM (4x) DebateGPT Actor Critic ACC-Debate ACC-Debate+ BoolQ .$798_{\pm.004}.$$848_{\pm.002}.$$873_{\pm.002}.$$843_{\pm.002}.$$877_{\pm.002}.$$\mathbf{.893}_{\pm.002}$ MMLU .$562_{\pm.005}.$$577_{\pm.002}.$$598_{\pm.002}.$$611_{\pm.001}.$$61_{\pm.005}.$$\mathbf{.672}_{\pm.004}$ BBH .$462_{\pm.003}.$$48_{\pm.012}.$$493_{\pm.012}.$$518_{\pm.005}.$$519_{\pm.009}.$$\mathbf{.601}_{\pm.004}$ SCIQ .$856_{\pm.002}.$$871_{\pm.003}.$$891_{\pm.001}.$$891_{\pm.002}.$$902_{\pm.005}.$$\mathbf{.905}_{\pm.002}$ ARC .$823_{\pm.002}.$$822_{\pm.002}.$$833_{\pm.003}.$$842_{\pm.003}.$$843_{\pm.003}.$$\mathbf{.856}_{\pm.003}$ Gemma-2 SoM (4x) DebateGPT Actor Critic ACC-Debate ACC-Debate+ BoolQ .$759_{\pm.004}.$$812_{\pm.003}.$$839_{\pm.005}.$$774_{\pm.014}.$$84_{\pm.005}.$$\mathbf{.845}_{\pm.005}$ MMLU .$578_{\pm.002}.$$\mathbf{.582}_{\pm.002}$$519_{\pm.026}.$$566_{\pm.002}.$$51_{\pm.016}.$$555_{\pm.003}.$ BBH .$449_{\pm.01}.$$491_{\pm.01}.$$51_{\pm.011}.$$475_{\pm.004}.$$\mathbf{.513}_{\pm.006}$$475_{\pm.008}.$ SCIQ .$903_{\pm.002}.$$914_{\pm.002}.$$\mathbf{.923}_{\pm.002}$$912_{\pm.001}.$$918_{\pm.003}.$$909_{\pm.003}.$ ARC .$843_{\pm.005}.$$851_{\pm.003}.$$85_{\pm.002}.$$\mathbf{.855}_{\pm.002}$$852_{\pm.003}.$$849_{\pm.002}.$
Table 2: Accuracy after 5 rounds of debate. The Actor (Critic) column corresponds to an actor-agent (critic-agent) trained via ACC-Debate and paired with an untrained Critic (Actor) during debate.
Figure 4: Comparison of responses from the critic model before and after training with ACC-Debate.
@@ 5.4 What do the Agents Actually Learn?
Lastly, we are interested in understanding how ACC-Debate improves the Actor-Critic Team. Figure 4 demonstrates an example of the difference in responses between an untrained critic and a critic trained through ACC-Debate. Although the actor provides a wrong answer, the untrained critic is too agreeable and does not provide substantive feedback for the actor to correct their answer.
@@ 6 Conclusion, Limitations and Impact
In this paper, we propose ACC-Debate, a novel framework for jointly training a two-agent team (one actor-agent and one critic-agent) to collaboratively solve problems through iterative discussion. To train these agents, we developed an off-policy data generation scheme dubbed “Guided-debate”, which produces high-quality preference data for collaborative models. We found that ACC-Debate outperforms all baselines on a wide array of domains. In particular, even a single round of training for both the actor and critic results in a high-quality debate team. Of particular note is the effects that ACC-Debate has on the critic model. Without ACC-Debate, the critic model often appears too agreeable in their responses. In contrast, after training with ACC-Debate, the critic is far more likely to provide detailed disagreements during debate. The success of ACC-Debate demonstrates the need for more specialized training when it comes to collaborative models.
However, our framework ACC-Debate also comes with limitations. Firstly, even though ACC-Debate attains superior performance compared to baselines on a wide array of domains, it is important to note that we conduct experiments mainly on question-answering tasks; thus, it remains to be seen whether such a framework would continue to be effective in other types of tasks. Our method makes use of the fact that for each question, correct and incorrect answers can be easily established. Secondly, while we provide results for three families of models, these experiments are performed on 2B, 7B, and 8B models. While our method is effective for these sizes (standard in open-source models), it remains to be seen whether this effectiveness will scale to larger models.
@@
Here, we outline the details necessary to reproduce our results. We provide an algorithm for our data generation procedure (Algorithm 1), as well as a description of our training procedure in Section 4.3. Each dataset, baseline method, and base model used are specified at the beginning of Section 5. We provide additional experimental details in Section A. Prompts used for our method can be found in Section C.
@@ Appendix A Experiments
Training for all models was performed via the trl library, using LoRAs of size 256. When training ACC-Debate with DPO we use a negative log-likelihood (NLL) regularization term (with weight $1$) as outlined in Pang et al. (2024a).
@@ Appendix B Preference Optimization
In the case of explicit reward maximization, the reward function $r(z^{(t)},x,y)$ can first be learned automatically by simply simulating debate. The most straightforward way to do this is to first simulate one full debate between the actor and critic, i.e., $\langle(z_{a}^{(0)},z_{c}^{(0)}),\ldots,(z_{a}^{(T)},z_{c}^{(T)})\rangle$. Then for each $z_{a}^{(t)}$, $z_{c}^{(t)}$ the remaining $T-t$ debate steps can be resampled to estimate the corresponding value of $r(z_{a}^{(t)},x,y)$ and $r(z_{c}^{(t)},x,y)$. These pairs, namely $\big{(}z_{a}^{(t)},r(z_{a}^{(t)},x,y)\big{)}$ and $\big{(}z_{c}^{(t)},r(z_{c}^{(t)},x,y)\big{)}$ can then be used to learn the reward function. This reward function can then be plugged into the desired preference optimization scheme.
Figure 5: Accuracy over five rounds of debate on BoolQ.
Figure 6: Accuracy over five rounds of debate on MMLU.
Figure 7: Accuracy over five rounds of debate on BBH.
Figure 8: Accuracy over five rounds of debate on SCIQ.
Figure 9: Accuracy over five rounds of debate on ARC.
@@ C.1 Prompts
Single-Shot Prompt (No Debate)
@@
Guided Single-Shot Prompt (No Debate)
@@
Debate Prompt for Actor
@@
Guided Debate Prompt for Actor
@@
Debate Prompt for Critic
@@
Guided Debate Prompt for Critic
prompt = (’I would like you to be a debate assistant. ’
@@ C.2 Critic Responses
Here, we present several examples of the critic’s responses when the actor provides the wrong answer; critic responses are shown before and after training with ACC-Debate. We observed a qualitative change in the critic’s responses after training with our pipeline; the critic was more willing to disagree with the actor, and the critic provided more detailed feedback.

### v3

@@ 5.3 Individual Performance
Next, we examine the relative effectiveness of both the actor and the critic. To do this, we train an actor and critic via ACC-Collab. Then, during deliberation, we pair the trained actor with an untrained critic and pair the trained critic with an untrained actor. In Table 2, column “Actor” corresponds to the former, while “Critic” corresponds to the latter. On average, the trained actor attains higher accuracy compared to the trained critic, this aligns with intuition as the actor is responsible for providing answers while the critic plays a supporting role. In most cases, the trained actor (paired with an untrained critic) outperforms SFT and DebateGPT. When the trained actor is paired with a trained critic (either ACC-Collab or ACC-Collab+), its performance is further improved.
Llama-3 SoM (4x) DebateGPT Actor Critic ACC-Collab ACC-Collab+ BoolQ .$811_{\pm.007}.$$815_{\pm.005}.$$867_{\pm.003}.$$834_{\pm.005}.$$887_{\pm.005}.$$\mathbf{.894}_{\pm.003}$ MMLU .$635_{\pm.004}.$$654_{\pm.005}.$$651_{\pm.012}.$$65_{\pm.008}.$$644_{\pm.01}.$$\mathbf{.683}_{\pm.012}$ BBH .$514_{\pm.005}.$$551_{\pm.008}.$$583_{\pm.01}.$$55_{\pm.015}.$$\mathbf{.593}_{\pm.006}$$574_{\pm.003}.$ SCIQ .$923_{\pm.002}.$$932_{\pm.001}.$$947_{\pm.002}.$$945_{\pm.001}.$$\mathbf{.952}_{\pm.0}$$948_{\pm.003}.$ ARC .$874_{\pm.001}.$$876_{\pm.002}.$$\mathbf{.885}_{\pm.003}$$866_{\pm.002}.$$881_{\pm.004}.$$869_{\pm.002}.$ Mistral SoM (4x) DebateGPT Actor Critic ACC-Collab ACC-Collab+ BoolQ .$798_{\pm.004}.$$848_{\pm.002}.$$873_{\pm.002}.$$843_{\pm.002}.$$877_{\pm.002}.$$\mathbf{.893}_{\pm.002}$ MMLU .$562_{\pm.005}.$$577_{\pm.002}.$$598_{\pm.002}.$$611_{\pm.001}.$$61_{\pm.005}.$$\mathbf{.672}_{\pm.004}$ BBH .$462_{\pm.003}.$$48_{\pm.012}.$$493_{\pm.012}.$$518_{\pm.005}.$$519_{\pm.009}.$$\mathbf{.601}_{\pm.004}$ SCIQ .$856_{\pm.002}.$$871_{\pm.003}.$$891_{\pm.001}.$$891_{\pm.002}.$$902_{\pm.005}.$$\mathbf{.905}_{\pm.002}$ ARC .$823_{\pm.002}.$$822_{\pm.002}.$$833_{\pm.003}.$$842_{\pm.003}.$$843_{\pm.003}.$$\mathbf{.856}_{\pm.003}$ Gemma-2 SoM (4x) DebateGPT Actor Critic ACC-Collab ACC-Collab+ BoolQ .$759_{\pm.004}.$$812_{\pm.003}.$$839_{\pm.005}.$$774_{\pm.014}.$$84_{\pm.005}.$$\mathbf{.845}_{\pm.005}$ MMLU .$578_{\pm.002}.$$\mathbf{.582}_{\pm.002}$$519_{\pm.026}.$$566_{\pm.002}.$$51_{\pm.016}.$$555_{\pm.003}.$ BBH .$449_{\pm.01}.$$491_{\pm.01}.$$51_{\pm.011}.$$475_{\pm.004}.$$\mathbf{.513}_{\pm.006}$$475_{\pm.008}.$ SCIQ .$903_{\pm.002}.$$914_{\pm.002}.$$\mathbf{.923}_{\pm.002}$$912_{\pm.001}.$$918_{\pm.003}.$$909_{\pm.003}.$ ARC .$843_{\pm.005}.$$851_{\pm.003}.$$85_{\pm.002}.$$\mathbf{.855}_{\pm.002}$$852_{\pm.003}.$$849_{\pm.002}.$
Table 2: Accuracy after 5 rounds of deliberation. The Actor (Critic) column corresponds to an actor-agent (critic-agent) trained via ACC-Collab and paired with an untrained Critic (Actor) during deliberation.
Figure 4: Comparison of responses from the critic model before and after training with ACC-Collab.
@@ 5.4 What do the Agents Actually Learn?
Lastly, we are interested in understanding how ACC-Collab improves the Actor-Critic Team. Figure 4 demonstrates an example of the difference in responses between an untrained critic and a critic trained through ACC-Collab. Although the actor provides a wrong answer, the untrained critic is too agreeable and does not provide substantive feedback for the actor to correct their answer.
@@ 6 Conclusion, Limitations and Impact
In this paper, we propose ACC-Collab, a novel framework for jointly training a two-agent team (one actor-agent and one critic-agent) to collaboratively solve problems through iterative discussion. To train these agents, we developed an off-policy data generation scheme dubbed “Guided-Collaboration”, which produces high-quality preference data for collaborative models. We found that ACC-Collab outperforms all baselines on a wide array of domains. In particular, even a single round of training for both the actor and critic results in a high-quality team. Of particular note is the effects that ACC-Collab has on the critic model. Without ACC-Collab, the critic model is often too agreeable and lacks verbosity in their responses. In contrast, after training with ACC-Collab, the critic is far more likely to provide detailed disagreements during discussion.
However, our framework ACC-Collab also comes with limitations. Firstly, even though ACC-Collab attains superior performance compared to baselines on a wide array of domains, it is important to note that we conduct experiments mainly on question-answering tasks; thus, it remains to be seen whether such a framework would continue to be effective in other types of tasks. Moreover, in our experiments, we train and test models on the same task (partitioning each task into a training and testing set). As such, the generalizability of each actor-critic team to unseen domains is unknown. Our method makes use of the fact that for each question, correct and incorrect answers can be easily established. Secondly, while we provide results for three families of models, these experiments are performed on 2B, 7B, and 8B models. While our method is effective for these sizes (standard in open-source models), it remains to be seen whether this effectiveness will scale to larger models.
Acknowledgments
We would like to thank Li Hang for his valuable insights and guidance during the development of this work.
@@ Acknowledgments
Here, we outline the details necessary to reproduce our results. We provide an algorithm for our data generation procedure (Algorithm 1), as well as a description of our training procedure in Section 4.3. Each dataset, baseline method, and base model used are specified at the beginning of Section 5. We provide additional experimental details in Section A. Prompts used for our method can be found in Section C. Lastly, we publicly release the code used for our method.
@@ Appendix A Experiments
Training for all models was performed via the trl library, using LoRAs of size 256. When training ACC-Collab with DPO we use a negative log-likelihood (NLL) regularization term (with weight $1$) as outlined in Pang et al. (2024a).
Figure 5: ACC-Collab training pipeline, exemplified for the actor (top) and critic (bottom). The process 1) We generate data from both natural deliberation as well as guided deliberation towards and away from the ground truth answer $y$ using the actor and critic. 2) We compute the relative quality of each trajectory based on the expected quality difference $\Delta_{y},\Delta_{!y}$ w.r.t. to the natural response. 3) We store all high-quality pairwise data in our database and train the actor model. 4) We alternate this procedure for the actor and critic. As outline in Section 4 both the guidance procedure and the computation of $\Delta_{y},\Delta_{!y}$ differ between the actor and critic.
A.1 Correctness of Reasoning Steps
In addition to measuring outcome accuracy (i.e., the accuracy of the actor’s final answer), we also measure process accuracy (i.e., the accuracy of the reasoning and discussion steps taken by the actor and the critic model). These results are displayed in Table 3. We observe that our training pipeline typically maintains or increases the accuracy of the reasoning steps.
Unlike outcome accuracy, process accuracy does not have ground truth answers. Thus we utilize GPT-4o to serve as an oracle. We use the following steps to evaluate the correctness of the reasoning and discussion steps.
  1. 1.
    For both an untrained actor-critic team and a trained actor-critic team (trained via ACC-Collab), we randomly sample 500 deliberations, which resulted in the correct answer after 5 rounds of deliberation.
  2. 2.
    We then prompted GPT-4o to evaluate the correctness of the reasoning and discussion steps taken by the actor-critic teams in each of the 500 deliberations.
  3. 3.
    We then compute the average accuracy of the reasoning and discussion steps as evaluated by GPT-4o.
Dataset ACC-Collab+ Accuracy ACC-Collab Accuracy Untrained Accuracy
Llama-3
BoolQ .$890_{\pm.035}.$$894_{\pm.034}.$$794_{\pm.045}.$
MMLU .$968_{\pm.020}.$$976_{\pm.017}.$$909_{\pm.032}.$
BBH .$519_{\pm.064}.$$575_{\pm.055}.$$610_{\pm.055}.$
SCIQ .$958_{\pm.022}.$$981_{\pm.015}.$$937_{\pm.027}.$
ARC .$882_{\pm.046}.$$833_{\pm.042}.$$883_{\pm.036}.$
Mistral
BoolQ .$907_{\pm.035}.$$928_{\pm.029}.$$894_{\pm.034}.$
MMLU .$882_{\pm.042}.$$884_{\pm.036}.$$878_{\pm.037}.$
BBH .$766_{\pm.045}.$$818_{\pm.043}.$$538_{\pm.056}.$
SCIQ .$973_{\pm.018}.$$972_{\pm.018}.$$933_{\pm.028}.$
ARC .$927_{\pm.029}.$$952_{\pm.024}.$$852_{\pm.040}.$
Gemma-2
BoolQ .$869_{\pm.038}.$$872_{\pm.037}.$$861_{\pm.039}.$
MMLU .$844_{\pm.041}.$$843_{\pm.041}.$$800_{\pm.045}.$
BBH .$597_{\pm.056}.$$637_{\pm.054}.$$514_{\pm.056}.$
SCIQ .$984_{\pm.014}.$$978_{\pm.016}.$$985_{\pm.014}.$
ARC .$850_{\pm.040}.$$854_{\pm.039}.$$834_{\pm.042}.$
Table 3: Accuracy of the actor’s and critic’s reasoning and discussion steps as evaluated by GPT-4o.
@@ Appendix B Preference Optimization
In the case of explicit reward maximization, the reward function $r(z^{(t)},x,y)$ can first be learned automatically by simply simulating deliberation. The most straightforward way to do this is to first simulate one full deliberation between the actor and critic, i.e., $\langle(z_{a}^{(0)},z_{c}^{(0)}),\ldots,(z_{a}^{(T)},z_{c}^{(T)})\rangle$. Then for each $z_{a}^{(t)}$, $z_{c}^{(t)}$ the remaining $T-t$ deliberation steps can be resampled to estimate the corresponding value of $r(z_{a}^{(t)},x,y)$ and $r(z_{c}^{(t)},x,y)$. These pairs, namely $\big{(}z_{a}^{(t)},r(z_{a}^{(t)},x,y)\big{)}$ and $\big{(}z_{c}^{(t)},r(z_{c}^{(t)},x,y)\big{)}$ can then be used to learn the reward function. This reward function can then be plugged into the desired preference optimization scheme.
Figure 6: Accuracy over five rounds of deliberation on BoolQ.
Figure 7: Accuracy over five rounds of deliberation on MMLU.
Figure 8: Accuracy over five rounds of deliberation on BBH.
Figure 9: Accuracy over five rounds of deliberation on SCIQ.
Figure 10: Accuracy over five rounds of deliberation on ARC.
@@ C.1 Prompts
Single-Shot Prompt (No Deliberation)
@@
Guided Single-Shot Prompt (No Deliberation)
@@
Deliberation Prompt for Actor
@@
Guided Deliberation Prompt for Actor
@@
Deliberation Prompt for Critic
@@
Guided Deliberation Prompt for Critic
prompt = (’I would like you to be a deliberation assistant. ’
@@ C.2 Critic Responses
Here, we present several examples of the critic’s responses when the actor provides the wrong answer; critic responses are shown before and after training with ACC-Collab. We observed a qualitative change in the critic’s responses after training with our pipeline; the critic was more willing to disagree with the actor, and the critic provided more detailed feedback.